OpenStudy (anonymous):

If a snowball melts so that the surface area S=4*pi*r^2 decreases at a rate of 3 cm^2/min, find the rate at which the radius decreases when the radius is 5cm?

OpenStudy (ash2326):

We have \[S= 4\pi r^2\] It's given that the snow ball decreases at a rate of 3 cm^2/min so \[\frac{ds}{dt}=-3 cm^2/min\] We have to find the rate of decrease of radius \(\frac{dr}{dt}\) when radius= 5 cm Let's find \(\frac{ds}{dt}\) \[\frac{ds}{dt}= 8 \pi r \frac{dr}{dt}\] we have \(\frac{ds}{dt}=-3\) and radius = 5cm \[ -3 = 8\pi \times 5 \frac{dr}{dt}\] we'll get \[\frac{dr}{dt}=-\frac{3}{40\pi}\] We get \[\frac{dr}{dt}=-0.02388 cm/minute\]

OpenStudy (anonymous):

Thank you

OpenStudy (ash2326):

welcome , did you understand it ?

OpenStudy (anonymous):

yeah i got down to the part of 8pi *40 but i wasnt sure that the 3 needed to be negative and i wasnt understanding how it was on top but i understand now why its that way

OpenStudy (ash2326):

great, since it's decreasing it's negative